OpenStudy (lgbasallote):

Let P(x,y) denote the statement "x=y+3". What are the truth values of the propositions P(1,2) and P(3,0)?

OpenStudy (unklerhaukus):

\[\begin{array}{|c|c|c|c|}\hline x&y&P(x,y)&T/F \\\hline 1&2&1=2+3 &F\\3&0&3=0+3&T\\\hline\end{array}\]

OpenStudy (lgbasallote):

how do you make a table in latex?

OpenStudy (unklerhaukus):

\[\text{\[\begin{array}{|c|c|c|c|}\hline x&y&P(x,y)&T/F}\]\[\text{ \\\hline 1&2&1=2+3 &F\\3&0&3=0+3&T\\\hline\end{array}\]}\]

OpenStudy (lgbasallote):

oh. hline. that's what i was missing

OpenStudy (lgbasallote):

anway...what did you do?

OpenStudy (unklerhaukus):

i just made table entries that corresponded to the points of proposition, and then found the statement of P at each point and lastly i checked if the statements were mathematically sound

OpenStudy (lgbasallote):

so it's just substitution?

OpenStudy (lgbasallote):

hmm son of a gun...you're right

OpenStudy (unklerhaukus):

i have never seen a question like this before, so their could be another way , but the method i have used (made up) seams to make perfect sense to me

OpenStudy (lgbasallote):

that was right actually

OpenStudy (unklerhaukus):

yay